Dome 430 MHz calibration
There is a 430 MHz receiver on the
gregorian dome. It can be used for transmitting and receiving. It
has native circular polarization.
Measurements in date order
tsys polB jumping around day to day. Problem is polb jennings
06jun18: highTsys polA bad connection on output of waveguide filter.
For aero run moved polBFilter to polA.
13sep13: ripple in polA (it's been there
awhile). Tapping on polA again caused jumps.
09sep13: polA cavity filter repaired (loose solder joint) and then
reinstalled. both jennings relays now look ok.
cavity filter (before amp) was removed. Its loss was
changing by up to 5db when tapped. This was probably causing the
polA Tsys jumps.
05sep13: polA jennings relay switched with
polB jennings relay (during radar run).
27aug13: polA jennings relay found to have
large resistance. Replaced with spare Jennings relay. Tsys back down
to around 60K
27aug13: damaged "new" turnstile replaced by
"old" 2006 turnstile. Tsys polA back to 60K.
05aug13: waveguide assembly damaged during ao19 removal. floor
rotated, and hook on hoist caught turnstile on the adjustable
04apr13: Today found that polB jennings relay had been installed
reversed. It was switching sky between load and lna, not lna between
sky and load. reinstalled correctly.
??mar13: jennings relay polB reinstalled. Back from repair.
14jan13: polA cable dewar to if/lo rack had problem .. it was
21jan13: polB amp failed (450-12) . probably the ch xmter since no
jennings relay on polB. replaced with repaired amp (450-18) 25jan13.
13jan13: updated pointing offsets.
03dev12: put spare cavity filter from chPolA in front of dome polA
to see if ripples went away (no).
30nov12: polA cavity filter removed to see if it caused the ripples
(no).Also check for digital tv intermods without filter (yes)
jennings relay polB bad. removed. 19oct12 reinstalled for some tests
and then removed again.
30jun11: radioastronomy,radar switch polA was bad. removed 27jun11
and repaired it. Probably causing the ripples in polA band pass
19jul06: tsys polB increased. Stayed high till about 10aug06. The
jennings relay for polB was bad.
18jul06: new turnstile installed. Better isolation. Still needs to
be tuned a bit.
25may06: power supply for cal driver circuitry was bad. replaced it.
23may06: tsys starting to go infinite.. turns out power supply for
cal driver went bad.
27jan05: filters (both pols) installed before dewar to get rid of tv
oct04: filter placed in front of dome polA dewar removes
Measurements in date
- 29aug19: Tsys change as
aeronomy receiver goes from cooled receiver to warm receiver.
polB jennings relay give high tsys.
- 05feb14: ratio of backscatter gain to
on axis gain.
- 13sep13: ripple seen in polA bandpass.
Tapping on plumbing causes polA jumps (again).
- 09sep13: polA cavity filter repaired an reinstalled.
- 05sep13: polA cavity filter (before
amp) bad. removed. Was causing jumps in tsys polA
- 05sep13: polA jennings relay bad.
switched with polB relay (at least for radar run).
- 27aug13: new turnstile
damaged. replaced by old turnstile. Jennings relay polA also
- 05aug13: "new" turnstile
damaged. replace by old turnstile 27aug13.
- 03apr13: polA Tsys jumps to 400 K in the
calibration results: jun12-dec12. updated pointing offset.
- 02dec12: sweep notched filter in front of
az swings with polA filter removed tv station intermods,az
dependence, time periodic birdies
a2704 total power mapping shows polA ripples for > 1/2 the
spectra while moving radar,radio astronomy switch. Ripple in
x111 data shows polA ripple coming and going.
polB jennings relay removed. Little power from sky was getting
- 15oct12: 430 dome polB gain
is 10% of the expected value.
- 19sep12: TsysPol A is jumping
around. Also plot Tsys vs Freq.
atm power profiles show cal dropouts.
terminate different positions before dewar looking at polA
- 26mar12 -
looked at polA ripple while on load.
polA ripple continues ..
ripple remains when radar/radioastronomy switch bypassed.
proj a2599 dps sees polA bandpass irregularities.
polA shows irregularities in its bandpass
- 2011 - 2012 ripples
- 24oct09: polB had normal
gain. So polB jennings relay ok here.
- 27sep10: check 430
stability.. looks ok. dynamic spectra, avg spectra, rms/mean,
total power vs time.
- 10aug09: polA unstable
- 23apr09: Noise at 439 MHz in
dome system when Tx Hv on.
spectrum at dewar output during az swing.
new turnstile (.jpeg)
old turnstile (.jpeg)
- 11jan04: (TsysPolA -
TsysPolB) drift for 2003, 2004
14,22 causing intermods in both ch and dome dewars.
190829 430 warms up during aeronomy run.
We were running an aeronomy experiment (using
the sterling cooled amp) on 29aug19 when the system warmed up.
I computed the Tsys for the start of each ri file using 10secs of
cal measurements for each point.
The plot shows the change in Tsys
from cold to warm receiver (.ps) (.pdf)
- Tsys went from about 180 K cold to 200K warm.
- The bump at 21 hours is the galaxy passing overhead (19:00
05feb14 ratio of backscatter gain to on
Herb carlson and frank djuth wanted the
backscatter gain for the 430 dome.
15 heiles calibration
patterns on B0316+162 (CTA21.. point source with a flux of 8.7 Jy)
were used. Carl's patterns consist of:
- 4 passes through the source. Each pass is 6 fwhm beams long
- az strip through source
- za strip through source.
- -45 degree strip through source
- +45 degree strip through source.
- A 2-d gaussian with sidelobes is fit to each pattern. The fit
- offset, peak, widths in 2 directions, rotation angle for
ellipse, coma parameter
- the sidelobes are fit as gaussians. an 8th order harmonic
fit is done at the sidelobe distance
- Each pattern had 4 separate measurements:
- 25Mhz bw centered at 430 Mhz
- 12.5 Mhz bw centered at 430Mhz
- 1.56 Mhz bw centered at 429 Mhz
- 1.56 Mhz bw centered at 431 Mhz
- The variations are to check for interference.
Computing the back scatter gain:
For each pattern (15*4=60) , a 200x200 pixel beam map (bmap) was
created by evaluating the fit values that pattern
The first image shows an
example of a single beam map (.gif)
Each map was normalized to unity (the on axis gain = 1).
- The image covers +/-35.4 arc minutes
- The contours are spaced by 3db.
- This pattern was taken at za = 16.8 deg.
the ratio of (integral(bmap^2)/integral(bmap) was then
computed. Multiplying this ratio by the on axis gain should give the
back scatter gain.
The plots show the results of the
computation (.ps) (.pdf)
- Page 1:
- top: az,za for the 15 patterns . The left side has the
source rising, the right have is the source setting.
- Bottom: Point source gain vs za for the measurements.
- the colors show the 4 bands taken for each pattern
- The 25Mhz band was lower than the other 3. this probably
includes the gain dropoff with freq.
- Page 2: over plot the central strips for each pattern
- Black is the azimuth strip through each map.
- red is the za strip through each map.
- the azimuth width is narrower because of the elliptical
illumination of the primary (extended in the az direction).
- Page 3: ratio of integral(bmap^2)/integral(bmap)
- Top all patterns. average value: .423
- Bottom: limit patterns to za > 10 (the dome sits above
za=10 for most world day runs). Average value .420.
- The integral was done over a radius of 35 arc minutes from
the beam center
13sep13: polA ripple
On 13sep13 we tested polA of 430 dome because
the tsys had started to jump around again. The test was:
spectral line data was also taken while tapping.
- On the oscilloscope we could see jumps in the total power when
denis lightly tapped on different parts of the plumbing between
the turnstile and the amp. He tapped on : the jennings relay, ra
switch, elbows, filter, cable to dewar.
- The jumps seemed to be stronger closer to the filter. They
were up to Tsys (while on sky).
- When the jenning relay was switched to load, tapping caused
jumps that were about the same size as on sky (even though
Tsys had increased by a factor of 5).
x111 data from 01aug13 -> 13sep13 was then
used to see how the bandpass rippled changed with different
configurations of the system.
- 25 Mhz, 1 sec dumps, center at 430 Mhz with the interim
correlator using 9 level sampling
- The spectra showed bandpass ripple in pola.
- The ripple was always present (with and without the
The plots show
the bandpass shapes (.ps) (.pdf):
- Page 1: 1 second spectra taken 13aug13.
- Black is polA, red is polB
- Top: spectra on sky
- Bottom: Spectra on load (jennings relay closed) and sky.
- upper spectra on load shows no ripple
- bottom spectra is the spectra from the top frame plotted
- The ripples go away when on load.
- Page 2,3 - overplot x111 spectra 01aug13 -> 13aug13
- Page 2 is polA
- Page 3 is polB
- The spectra are plotted with offsets (for display). 01aug is
at the bottom,
- Colors show different configurations for the 430:
|polA filter bad
|the ripples are gone
- PolA has had ripple in the bandpass since at least aug13. PolB
does not have this ripple
- broken wires were found in the polA cavity filter on 05sep13.
- In the lab, tapping the filter with the wire broken caused
large jumps in the filter transmission.
- It was fixed and replaced on 09sep13.
- In the lab, tapping on the filter did not change the
transmission properties (after the fix).
- The ripple was present before the turnstile was damaged.
- The turnstile damage did not make the ripple worse.
- putting in the old turnstile seems to have increased the
- Using the radar receiver without the filter made the ripples
- This means that the problem is not the jennings relay
- repairing the filter did not fix the ripple problem.
- tapping on the plumbing after the filter fix caused jumps in
the power after installation in the dome (although this was not
checked until 13sep13).
05sep13: polA cavity filter removed. polA
jennings replaced. (top)
On 05sep13 around 10:00 the aeronomy
experiment was stopped to switch the radar room temp miteq
amp with a cooled version.
- Prior to the replacement we were getting tsys around 175K
- after the replacement we got Tsys values that varied from 105
- Switching from load to antenna and back caused the antenna
level to jump by up to 3db.
- After some more tests it was decided that the polA jennings
relay (installed 27aug13) was not working.
- A spare relay driver was tested, but the problem remained.
- The jennings relay on polB was switched with the polA
jennings relay. during radar runs, polB is not used.
- We measured a tsys of about 150K after the jennings relay
- After the switch we was saw large jumps in tsys when the
plumbing between the turnstile and the dewar was touched/banged.
The jumps were especially strong when areas near the cavity
filter were hit.
- We removed the cavity filter. the miteq amps have more dynamic
range than the cooled amps.
- After removing the cavity filter:
- Tsys was 95K in polA
- Tsys did not jump when we banged on parts of the plumbing.
- The filter was brought down to the lab and checked:
- on the network analyzer, lightly tapping the filter
or connectors caused level jumps of up to 5 db.
- This filter is probably why tsys PolA had been jumping
around (for many months).
- A loose solder joint was found inside the cavity filter. It
was fixed and then reinstalled in the dome 09sep13.
- the jennings relay on polB (originally from polA ) was
exercised (after the radar run). No problems were found. The
problem must have been that the polA relay was shaking the
filter a little more than the polB. The jennings relays
05aug13: turnstile damaged (top)
The turnstile (install 2006) was damaged on
05aug13 while removing ao19. The winch was left hanging and hooked
the adjustable backshort on the edge of the turnstile when
the rotary floor was moved. This bent the turnstile.
Tsys polA remained 200-300 K till 16aug13..
- when it occurred:
- 05aug13 7:12 am. tsys 80.7,79.3 K
- 05aug13 11:24 am. tsys: 323.3, 60.4 K
The damaged turnstile was replaced by the pre 2006 turnstile on
- the "old" turn stile does not have a tuning stub. it's
centerfreq is lower than 430 Mhz.
- Tsys polA had jumped up to 250-300 K after the problem
The jennings relay polA was found to have large impedance.. It
was also replaced on 27aug13 when the old turnstile was
installed. Tsys went back down to around 60K.
03apr13: polA Tsys jumps high in
04apr13: resolution: polB
jennings relay had been installed backwards. It was switching
sky between load and lna instead of lna between load and sky.
reinstalled correctly. This means that the high polA Tsys
was probably both were on load.
Since around the 31mar13 some afternoon tsys
measurements have shown TsysPolA to be much higher than polB (250
to 400 K vs 100)
Plots were made of the hcorcal tsys for 430 using data 21mar13
The plots show tsys vs date, hour,
and lst (.ps) (.pdf):
- PolA is black, polB is red.
- Top: 430 dome Tsys vs date .. starting at 21mar13 thru 03apr13
- The two largest polA tsys' occurred on 31mar13,and 01apr13
- middle: Tsys vs hour of day (AST)
- The large Tsys occurred at 15:30 and 16:00 hours
- Bottom: Tsys vs local sidereal time.
- Tsys will increase at 6:00 (galactic anti center) and 19:00
- PolA,B will not differ this much because of polarized power
from the sky..
- The two high measurements have nearly identical values even
though they are separated by 24 hours.
- if this was rfi i would expect the differences to jump
- I think the high value is close to Tsys with the receiver on
- My guess this is a problem with the load switch for polA.
- If the operator accidentally left the load switch on load,
then polB should also be high (need to check polB load)
- There may be problem with the load switch or the
signal going to it.
02dec12: sweep notched filter in front of
A single pole cavity filter (with notches for the
analog tv stations) was placed in front of the dewar 27jan05. The
polA filter had problems with the input connector (a c-clamp was
added to replace one of the 4 screws holding the connector). This
filter was replaced by one of the single pole cavity filters from
the carriage house (around 02dec12) The filter with the
connector problem was swept by dana (down in the lab).
plots show the filter response (.ps) (.pdf):
- the filter was swept 3 times.
- Each had a differing tension on the filter connector
to see how much it would change.
- black,red,green show the 3 different measurements.
- The 4 frames show S11, S22, S21, S12
- S11 is the reflection from the damaged connector. This was
connected to the antenna
- S22 is the reflection from the dewar side of the filter
- S21 is the transmission from the antenna to the dewar
- S12 is the transmission from the dewar to the antenna
- Page 1:
- The entire freq 180 to 680 MHz
- Page 2: blowup showing 400-450 MHz
- Page 3: blowup 460 to 530 MHz
- The dashed vertical lines are the tv station bands.
- Chn 14 and 22 were the analog stations causing the intermods
in the dewar.
- With the switch to digital tv, these stations are no
longer very strong (22 in not present at all).
- Chn 15 and 23 (digital stations) are now causing intermods
in the dewar.
- Then notch near chn 15 changed a lot when for the sweeps
with different amounts of stress.
- chn 23 is not cut off very much by the old notches.
27sep10: check 430
PolA 430gr was unstable back on 10aug09. On
27sep10 data was taken to see if the problem still persisted. The
- Telescope sitting at az=285, za=9.
- mock spectrometer took 21 MHz centered at 430MHz. The 422-442
filter was in.
- 300 seconds of .1 sec averages were taken.
data images, plots:
- There are no large jumps or drifts in the total power vs time
for polA, or polB (unlike aug09).
- the 1.2 hz spike in polA total power spectra is gone
10aug09: 430 polA unstable (top)
Project a2415 reported that 430 dome polA was
unstable with a 15 second periodicity. Data was taken on 10aug09 AM
with the mock spectrometers. The setup was:
The total power for the 300 seconds of data was
computed for each .1 second sample after throwing out rfi (compute
rms/mean by channel and throw out channels with large rms). The
total power was then normalized to the median value over the 300
- az,za=270,1.09 sitting.
- mock spectrometer: 8k channels, 34 MHz Bw, and .1 second
- Data taken for 300 seconds.
The plots show the total power and
spectrum of the total power (.ps) (.pdf) :
- Black is polA, red is polB
- Top: total power vs time
- Middle: The spectrum of the total power was computed and
plotted. The 1.2 hz and harmonics come from the dewar
- Bottom: blowup of the spectrum of the total power (0-1 hz).
There is no obvious 15 second periodicity (.066 hz). Maybe what
people saw is not very stable in time.
- Pol A has a large 1.2 hz periodicity in its total power. This
is coming from the refrigerator. We should probably bang on the
can to see if something is loose inside...
mar03 thru oct03 GAIN CURVE. (top)
Calibration data taken mar03 thru oct03 was used
to compute a gain curves. The fits were done versus za. The plot
the data, fits, and the residuals. This needs to be redone
once the filters are installed on the frontend (to cut out the
- TsysPolB drift for 2003 and 2004 (top)
The system temperature for the dome 430 receiver is
measured regularly during the normal maintenance
procedure. I've plotted Tsys
years 2003 and 2004 (.ps). (.pdf)
Fig1- Top: Tsys for 2003. Black is
polA, red is polB, Bottom: TsyspolA-tsysPolB
Fig2:- Top: Tsys for 2004. Bottom:
From the plots you can see that the difference
TsysPolA-TsysPolb has been changing. In july 2003 the difference
started to increase (TsysA larger then TsysB). Around 17nov03 the
difference jumped by 10 degK so that TsysB was larger then tsysA
by about 10 degK. It stayed at this level until august
2004 when it started to increase gradually.
The tsys measurements tend to be
taken at 16:30 AST so the RA of the measurement will move during
the year. It seems unlikely that the gradual drift in
TsysPolA - TsysPolB is coming from the sky since the receiver is
circularly polarized (and there are no extended sources of
circular polarization that i know of). Looking at the
src deflections over this time period the polA,polB ratio remained
fairly constant, so my guess is that Tsys was changing (and not
the cal size).